OpenStudy (lgbasallote):

obtain a solution using linear DE from y' = exp (2x - y)

OpenStudy (anonymous):

exp?

OpenStudy (anonymous):

its in variable separable form...... \[\frac{dy}{dx}=\frac{e ^{2x}}{e ^{y}}\] \[dy(e ^{y})=dx(e^{2x})\]

OpenStudy (lgbasallote):

but how to do it with linear DE only?

OpenStudy (anonymous):

linear DE should be of the form \[A \frac{d^{n}y}{dx^{n}}............+B \frac{d^{2}y}{dx^{2}}+Cfrac{dy}{dx}=P\]

OpenStudy (anonymous):

we dont get this form here

OpenStudy (lgbasallote):

soit's impossible?

OpenStudy (anonymous):

i guess so ...not sure..

OpenStudy (lgbasallote):

what about dy/dx - e^2x/e^y = 0?

OpenStudy (anonymous):

dats not linear....

OpenStudy (lgbasallote):

oh hmm so it's really impossible?

OpenStudy (anonymous):

guess so!!!